Q.1 ARE THE FUNCTIONS EVEN OR ODD

1- INTEGRAL IS FROM MINUS INFINITY TO PLUS INFINITY
∫e^(-2x)(1-e^(-x))^(2) dx

2-INTEGRALIS FROM MINUS INFINITY TO PLUS INFINITY
∫x(e^(-2x))(1-e^(-x))^(2) dx
1
Expert's answer
2020-04-17T17:57:51-0400

1.

"f(x)=_{-\infty}^{\infty}\int e^{-2x}(1-e^{-x})^2dx"

Changing x with -x we get;

"f(-x)=_{-\infty}^{\infty}\int e^{2x}(1-e^{x})^2dx"


Now "f(x)+f(-x) \neq 0"

And also "f(x)-f(-x) \neq 0"

Thus, f(x) is neither even nor odd.


2. Let "g(x)=_{-\infty}^{\infty}\int x e^{-2x}(1-e^{-x})^2dx"

Changing x with -x, we get;

"g(-x)=_{-\infty}^{\infty}\int- x e^{2x}(1-e^{x})^2dx"


Now, "g(x)+g(-x) \neq 0"

And also "g(x)-g(-x) \neq 0"

Thus, g(x) is also neither even nor odd.